# What Makes Good Synthetic Training Data for Zero-Shot Stereo Matching?

### v2

@@ 1 Introduction
Our study yields a variety of interesting findings. Of all the designs tested, we found the most effective stereo data comes from the combination of realistic indoor scenes with added floating objects. Removing realistically arranged background furniture (leaving floating objects in empty rooms) or removing the entire background (leaving just floating objects) harms zero-shot generalization. We also find that existing stereo networks struggle to learn reflective and transparent materials without large performance drops on diffuse regions. These results show that certain aspects of scene realism are helpful for zero-shot stereo, while also suggesting that scenes designed to be fully realistic lack sufficient dense geometry to be data-efficient.
@@ 2 Related Work
Non-procedural synthetic datasets [3, 21] such as IRS [31], TartanAir [32], UnrealStereo4k [29], and DynamicReplica [16] typically render frames from realistic scenes handcrafted by 3D artists. Simulators such as HS-VS [40] and VirtualKITTI [9] [4] primarily target driving scenarios.
Although these datasets have been widely adopted and have demonstrated success in state-of-the-art stereo networks, they do not report ablations over dataset generation parameters, and therefore provide limited insight into what actually makes these datasets so effective. Moreover, we cannot directly study generation parameters in existing datasets (e.g. re-rendering TartanAir with different materials) because they do not open-source their generation code or assets. We use the open-source Infinigen system to reproduce and analyze many existing design choices. We perform a detailed analysis of what parameters, such as object placement, camera parameters, and materials, matter most for downstream performance.
FoundationStereo [36] achieves state-of-the-art zero-shot performance by introducing a new model architecture alongside the FoundationStereo dataset (FSD). FSD adopts similar settings to our final dataset, such as combining floating objects with realistic background scenes and varying camera parameters. However, FoundationStereo does not provide any analysis of FSD or code to reproduce or expand its dataset. Our work provides this analysis and an open-source procedural generator to enable further research.
Figure 2: WMGStereo-150k Dataset. From left to right, we show random, non-cherrypicked samples from our Indoors with Floating Objects, Dense Floating Objects, and Nature scene types. Refer to the Supplement for additional samples.
@@
Our work significantly expands on the object and scene generators from Infinigen [23, 24]. We develop a procedural placement algorithm for floating object and lighting placement integrated with Infinigen scene and object generators. Our algorithm supports common features in stereo datasets and enables us to exercise fine-grained control over objects, materials, lighting, and cameras. For instance, we can remove different classes of objects from generation and even remove specific materials from individual objects. We open-source our configurations, enabling others to easily create targeted data, such as scenes with extreme clutter or all glass materials. Our dataset is a novel demonstration of general-purpose scene generators for stereo data.
Analysis of Synthetic Datasets Mayer et al. [20] perform a detailed study of the procedural generation parameters of synthetic training data for optical flow and disparity prediction. This study concludes that “realism is overrated” for effective synthetic data, motivating the further development of successful flow datasets such as AutoFlow [28, 13]. However, this study primarily studies generation parameters for FlyingChairs-style datasets [8, 14], which consist only of warped 2D images with no depth or disparity. Moreover, their study of disparity prediction focuses on augmentations simulating camera noise and blurring. As such, there is limited study of critical generation parameters, such as object placement and realism, for modern stereo datasets rendered from 3D scenes. Our study focuses on stereo depth prediction, uses more realistic 3D procedural generators, and applies directly to modern stereo networks. We find that a mix of placement realism and random floating objects is an effective recipe for stereo data. While this result supports Mayer et al.’s conclusion that diversity from randomly scattered objects is important, it also suggests that scene realism can help significantly with zero-shot generalization.
@@ 3.2 Parameter Study
We perform a systematic study to determine the effect of procedural generation parameters on zero-shot stereo matching performance. For a variety of parameters, we directly evaluate their effect on zero-shot performance. In each such case, we generate 5000 stereo pairs using the indoor floating objects scene type. We show a single illustrative example of each dataset in Fig. 1. We train RAFT-Stereo [18] on each of these datasets for 75k steps using standard training settings. Finally, we evaluate the resulting models on the Middlebury [26], ETH3D [27], KITTI [10, 22], and Booster [25] datasets and report results in Tab. 1. This training setup uses less data and a shorter training schedule to enable fast experimentation. In practice, we find that this reduced scale is sufficiently representative of the full-scale training runs.
We evaluate zero-shot performance on the Middlebury 2014 evaluation set, the Middlebury 2021 dataset, the ETH3D train set, the KITTI 2012 and KITTI 2015 train sets, and the Booster train set. For each dataset, px error is defined as the percentage of pixels with end-point-error exceeding a specified threshold. The thresholds are chosen using standard ranking metrics: 2px non-occluded error for Middlebury 2014, 2px error for Middlebury 2021, 1px error on ETH3D, 3px error for KITTI 2012 and KITTI 2015, and 2px error on Booster. We evaluate on full (F) and half (H) resolution on Middlebury and on quarter (Q) resolution on Booster.
@@
Figure 3: End-point-error averaged by object and material. We show average error only for the materials that make up at least 0.1% of pixels. Assets marked in red were removed from our system after manual inspection because they introduce ambiguous high-error cases, such as entirely transparent or reflective surfaces, or thin imperceptible foliage / holes.
@@ 3.3 Data Generation Cost Optimization
Generating data with the default Infinigen configurations is computationally expensive. The two largest costs are CPU time for generating the scene, and GPU time for rendering images from the generated scene, which we aim to reduce.
@@
WMGStereo-150k 5.10 3.76 6.72 2.50 3.30 4.54 9.09
FoundationStereo 5.80 3.27 6.93 2.13 3.56 4.18 7.51
Table 5: Zero-shot performance by dataset. DLNR trained only on WMGStereo-150k outperforms the same model trained on widely used stereo matching training datasets, and is competitive with training on the FoundationStereo dataset.
@@ 4 WMGStereo-150k Dataset
We use the best settings in Tab. 1 to construct WMGStereo-150k, a stereo dataset of 163,666 stereo pairs. This dataset incorporates all the best parameter configurations found in Sec. 3. We include three categories of rendered scenes, shown in Fig. 2: indoor floating objects (71,947 stereo pairs), dense floating objects (70,300 stereo pairs), and nature scenes (21,419 stereo pairs).
@@
We train RAFT-Stereo [18], DLNR [45], and Selective-IGEV [33] on our dataset for 200k steps, following the standard training procedures and hyperparameters used by the original authors. During training, we reweight the scene types on our dataset such that each type is sampled equally, following the results from the study in Section 3. During training, we mask the sky and the untextured room exteriors. We empirically found this to provide a small performance boost.
@@ 4.3 Dataset Comparison
We also compare our dataset individually to existing datasets. For each of SceneFlow, CREStereo, TartanAir, and IRS, we train DLNR for 200k steps using default training settings. We report quantitative zero-shot results in Tab. 5, using metrics as defined in Sec. 3. The model trained on WMGStereo-150k achieves better performance than classic datasets, with a 39% improvement over IRS on Middlebury 2014 (H), and is competitive with FSD. Despite not explicitly modeling driving scenarios like SceneFlow and FoundationStereo, our dataset achieves better performance than both datasets on KITTI-12. To demonstrate that our dataset has not overfit to the benchmarks in the study, we also evaluate on held out benchmarks not used in parameter tuning in Tab. 6. We use the Middlebury 2014 training set, excluding the standard 10 validation scenes, and a fixed, randomly selected subsample of 200 stereo pairs from the DrivingStereo [41] dataset.
Dataset Middlebury-Train DrivingStereo
@@
DLNR-WMGStereo-150k 12.61 4.54 1.89 0.82
Table 6: Zero-shot performance by dataset on held-out benchmarks Training DLNR on WMGStereo-150k results in better performance on benchmarks not used in parameter adjustment, demonstrating our findings generalize to unseen scenarios.
@@ 5 Conclusion
In this work, we contribute a thorough study on what matters in stereo dataset generation. By carefully selecting the best design choices, we generate a new synthetic dataset and demonstrate that training solely on this dataset achieves stronger zero-shot performance than training on many widely used datasets combined. We believe that our study and open-source codebase will be continually useful for stereo matching research, especially in co-developing training data with new architectures.
@@ Appendix A Additional Analysis
Appendix B Additional Generation Details
B.1 Indoor Floating Objects Scene Type
@@
B.2 Dense Floating Objects Scene Type
@@
B.3 Nature Scene Type
@@ B.4 Ground Truth
B.5 Dataset post-processing
Certain types of undesirable scenes are hard to prevent during dataset generation, so we provide tools to remove them as a post-processing step. We use a simple mean of pixel intensities to filter out frames that are extremely dark as a result of lighting augmentation. We remove all frames with any depth value lower than a certain threshold to eliminate scenarios where the camera clips into an object or when an object has extremely large disparity values. This threshold is set to 12.5 cm for indoor scenes and 5 m for nature scenes.
B.6 Atmospheric effects
@@
B.7 List of Assets Used

### v3

# arXiv:2504.16930v3[cs.CV] 27 Feb 2026

@@ 1 Introduction
Our study yields a variety of interesting findings. Of all the designs tested, we found the most effective stereo data comes from the combination of realistic indoor scenes with added floating objects. Removing realistically arranged background furniture (leaving floating objects in empty rooms) or removing the entire background (leaving just floating objects) harms zero-shot generalization. These results show that certain aspects of scene realism are helpful for zero-shot stereo, while also suggesting that scenes designed to be fully realistic lack sufficient geometry diversity to be data-efficient. We also find that existing networks struggle to learn reflective and transparent materials without large performance drops on diffuse regions, suggesting that co-design of architecture and data might be necessary for robust non-Lambertian stereo. Finally, we find that varying camera baseline greatly improves generalization, and that ray-tracing quality can be traded off for more data.
@@ 2 Related Work
Figure 2: WMGStereo-150k Dataset. From left to right, we show random, non-cherrypicked samples from our Indoors with Floating Objects, Dense Floating Objects, and Nature scene types. Refer to the appendix for additional samples.
@@
Non-procedural synthetic datasets [3, 24] such as IRS [35], TartanAir [36], TartanGround [26], XR-Stereo [6], DISC [17], UnrealStereo4k [33], and DynamicReplica [19] typically render frames from realistic scenes handcrafted by 3D artists. Simulators such as HS-VS [44] and VirtualKITTI [10] [4] primarily target driving scenarios.
Although these datasets have been widely adopted and have demonstrated success in state-of-the-art stereo networks, they do not report ablations over dataset generation parameters, and therefore provide limited insight into what actually makes these datasets so effective. Moreover, we cannot directly study generation parameters in existing datasets (e.g. re-rendering TartanAir with different materials) because they do not open-source their generation code or assets. We use the open-source Infinigen system to reproduce and analyze existing design choices. We perform a detailed analysis of what parameters, such as object placement and materials, matter most for downstream performance.
FoundationStereo [40] achieves state-of-the-art zero-shot performance by introducing a novel architecture alongside the FoundationStereo dataset (FSD). FSD simultaneously introduces many new dataset features, such as varied camera parameters, flying objects, randomized lighting, realistic background scenes, and physics-based simulation. Our parameter study disentangles and identifies which of these design choices are most critical for downstream performance, informing future dataset design. FSD is also a static dataset, while our dataset generation code is open-source, allowing our dataset to be expanded and tailored for target use cases. A user targeting non-Lambertian stereo could easily regenerate our data with only glass materials, while FSD provides limited coverage of transparent objects. Our dataset generator therefore uniquely enables the co-design of data with new tasks and model architectures.
@@
Our work significantly expands on the object and scene generators from Infinigen [27, 28]. We develop a procedural placement algorithm for floating object and lighting placement integrated with Infinigen scene and object generators. Our algorithm supports common features in stereo datasets and enables us to exercise fine-grained control over objects, materials, lighting, and cameras. For instance, we can remove different classes of objects from generation and even remove specific materials from parts of an individual objects. Our dataset is a novel demonstration of general-purpose scene generators for stereo data.
Analysis of Synthetic Datasets Mayer et al. [22] perform a detailed study of the procedural generation parameters of synthetic training data for optical flow and disparity prediction. This study concludes that “realism is overrated” for effective synthetic data, motivating the further development of successful flow datasets such as AutoFlow [32, 15]. However, this study primarily studies generation parameters for FlyingChairs-style datasets [9, 16], which consist only of warped 2D images with no depth or disparity. Moreover, their study of disparity prediction focuses on augmentations simulating camera noise and blurring. As such, there is limited study of critical generation parameters, such as object placement and realism, for modern stereo datasets rendered from 3D scenes. More recently, StereoCarla [14] studied the effects of camera placement and weather effects for driving simulator datasets.
In contrast to Mayer et al., our study focuses on stereo depth prediction, uses more realistic 3D procedural generators, and applies directly to modern stereo networks. We find that a mix of placement realism and random floating objects is an effective recipe for stereo data. While this result supports Mayer et al.’s conclusion that diversity from randomly scattered objects is important, it also suggests that scene realism can help significantly with zero-shot generalization. We also study both high-level design choices such as floating object placement and low-level parameters such as materials.
@@ 3.2 Parameter Study
We perform a systematic study to determine the effect of procedural generation parameters on zero-shot stereo matching performance. For a variety of parameters, we directly evaluate their effect on zero-shot performance. In each such case, we generate 5000 stereo pairs using the indoor floating objects scene type. We show a single illustrative example of each dataset in Fig. 1. We train RAFT-Stereo [21] on each of these datasets from a random initialization for 75k steps using standard hyperparameters and augmentation. Finally, we evaluate the resulting models on the Middlebury [30], ETH3D [31], KITTI [11, 25], and Booster [29] datasets and report results in Tab. 1. This training setup uses less data and a shorter training schedule to enable fast experimentation. In practice, this reduced scale is sufficiently representative of the full-scale training runs.
We evaluate zero-shot performance on the Middlebury 2014 evaluation set, the Middlebury 2021 dataset, the ETH3D train set, the KITTI 2012 and KITTI 2015 train sets, and the Booster train set. The px error is defined as the percentage of pixels with end-point-error exceeding a specified threshold. The thresholds are chosen using standard ranking metrics: 2px non-occluded error for Middlebury 2014, 2px error for Middlebury 2021, 1px error on ETH3D, 3px error for KITTI 2012 and KITTI 2015, and 2px error on Booster. We evaluate on full (F) and half (H) resolution on Middlebury and on quarter (Q) resolution on Booster.
@@
Figure 3: End-point-error averaged by object and material. We show average error only for the materials that make up at least 0.1% of pixels. Assets marked in red were removed after manual inspection because they introduce ambiguous high-error cases, such as completely reflective surfaces, or imperceptible holes.
@@ 3.3 Data Generation Cost Optimization
Generating data with the default Infinigen configurations is computationally expensive. We aim to reduce the two largest costs: CPU time for generating the scene, and GPU time for rendering images from the generated scene.
@@
FSD 5.80 3.27 6.93 2.13 3.56 4.18 7.51
[…]
Table 5: Zero-shot performance by dataset. DLNR trained only on WMGStereo-150k outperforms the same model trained on widely used datasets. Combining our data with FSD leads to best overall performance, demonstrating that our data can augment existing datasets.
@@ 4 WMGStereo-150k Dataset
We use the best settings in Tab. 1 to construct WMGStereo-150k, a stereo dataset of 163,396 stereo pairs. This dataset incorporates all the best parameter configurations found in Sec. 3. We include three categories of rendered scenes, shown in Fig. 2: indoor floating objects (71,800 stereo pairs), dense floating objects (70,300 stereo pairs), and nature scenes (21,296 stereo pairs).
@@
We train RAFT-Stereo [21], DLNR [49], and Selective-IGEV [37] on our dataset for 200k steps from a random initialization, following the standard training procedures and hyperparameters used by the original authors. During training, we reweight the scene types on our dataset to sample each one equally, following results from Section 3. During training, we mask the sky and untextured room exteriors.
@@ 4.3 Dataset Comparison
We also compare our dataset individually to existing datasets. For each of SceneFlow, CREStereo, TartanAir, and IRS, we train DLNR for 200k steps using default training settings. We report quantitative zero-shot results in Tab. 5, using metrics as defined in Sec. 3. The model trained on WMGStereo-150k achieves better performance than classic datasets, with a 39% improvement over IRS on Middlebury 2014 (H). Despite not explicitly modeling driving scenarios like FoundationStereo, our dataset achieves better performance on KITTI-12. Training on a combination of FSD and WMGStereo-150k results in better performance than either dataset alone on Middlebury (H), ETH3D, and Booster, showing our data is useful when mixed with other datasets.
To demonstrate that our dataset has not overfit to the benchmarks in the study, we evaluate on held out benchmarks not used in parameter tuning in Tab. 6. We use the Middlebury 2014 train set, excluding the 10 validation scenes, and a fixed, random subsample of 200 pairs from DrivingStereo [45]. Training on our data results in a 27% improvement over FSD in 3px error on DrivingStereo.
Dataset Midd-Train DrivingStereo
@@
[…]
Table 6: Zero-shot performance by dataset on held-out benchmarks Training on WMGStereo-150k results in better performance on benchmarks not used in parameter adjustment, demonstrating our findings generalize to unseen scenarios.
@@ 5 Conclusion
We contribute a thorough study on what matters in stereo dataset generation. By carefully selecting the best design choices, we generate a new synthetic dataset and demonstrate that training solely on this dataset achieves stronger zero-shot performance than training on many widely used datasets combined. We believe that our study and open-source codebase will be continually useful for stereo matching research, especially in co-developing training data with new architectures.
@@ 7 Additional Analysis
8 Additional Generation Details
8.1 Indoor Floating Objects Scene Type
@@
8.2 Dense Floating Objects Scene Type
@@
8.3 Nature Scene Type
@@ 8.4 Ground Truth
8.5 Dataset post-processing
Certain types of undesirable scenes are hard to prevent during dataset generation, so we provide tools to remove them as a post-processing step. We use a simple mean of pixel intensities to remove frames that are extremely dark as a result of lighting augmentation. We filter frames depth value lower than 12.5 cm to eliminate scenarios where the camera partially clips into an objects. We optionally apply a more aggressive distance filter of 5 m for the nature spit to filter out extremely high disparity samples, and release both the filtered and unfiltered nature split.
8.6 Atmospheric effects
@@
8.7 List of Assets Used